# Proximal Interacting Particle Langevin Algorithms

### v1

@@ Appendix B Theoretical Analysis of Proximal Particle Gradient Descent
where $\xi_{n}=B_{n\gamma}-B_{(n-1)\gamma}$ for $n\geq 0$ are $d_{x}$-dimensional standard Gaussians.
@@ B.2 Proximal Particle Gradient Descent
We extend the results of Caprio et al. [2024] to provide a non-asymptotic bound for MYPGD. To do so, we consider the following metric on $\mathbb{R}^{d_{\theta}}\times\mathcal{P}_{2}(\mathbb{R}^{d_{x}})$
@@ Appendix C Convergence to Wasserstein gradient flow
We start by proving the following auxiliary result. Let us denote, for any $\theta\in\mathbb{R}^{d_{\theta}}$ and $\nu\in{\mathcal{P}}(\mathbb{R}^{d_{x}})$, $g(\theta,\nu):=\int_{\mathbb{R}^{d_{x}}}\nabla_{\theta}U^{\lambda}(\theta,x^{\prime})\nu(x^{\prime})\mathrm{d}x^{\prime}$.
@@
The function $g:\mathbb{R}^{d_{\theta}}\times{\mathcal{P}}(\mathbb{R}^{d_{x}})\to\mathbb{R}^{d_{\theta}}$ is Lipschitz continuous in both arguments, i.e.,
@@
$\mathbb{E}\left[\sup_{t\in[0,T]}\left(\|\bm{\theta}_{\lambda,t}-\bm{\theta}^{N}_{t}\|+\|\bm{X}_{\lambda,t}-\bm{X}_{t}^{j,N}\|\right)\right]\leq\frac{\sqrt{2}(\sqrt{C_{T}}\lambda^{-1}+\sqrt{T})e^{2T\lambda^{-1}}}{N^{1/2}}$ (45)
where $C_{T}:=\sup_{t\leq T}\mathbb{E}\left[|\bm{\theta}^{N}_{t}|^{2}+|\bm{X}_{t}^{j,N}|^{2}\right]<\infty$, for any $j=1,\dots,N$.
@@
$\mathbb{E}\left[\sup_{t\in[0,T]}\left(\|\bm{\theta}_{\lambda,t}-\bm{\theta}_{t}\|^{2}+\|\bm{X}_{\lambda,t}-\bm{X}_{t}\|^{2}\right)\right]\leq\left(\lambda^{2}\|\nabla U\|_{\textnormal{Lip}}^{4}C_{T}+\|\nabla U\|_{\textnormal{Lip}}^{2}\mathcal{O}(\lambda^{4})\right)T\exp(2\|\nabla U\|_{\textnormal{Lip}}^{2}T),$
where $C_{T}:=\sup_{t\leq T}\mathbb{E}\left[|\bm{\theta}_{\lambda,t}|^{2}+|\bm{X}_{\lambda,t}|^{2}\right]<\infty$. It follows that, as $\lambda\to 0$, (44) converges to (46) in $\mathbb{L}^{2}$.
@@
$\bm{X}_{t}$ $=X_{0}-\int_{0}^{t}\nabla_{x}U(\bm{\theta}_{s},\bm{X}_{s})\mathrm{d}s+\sqrt{2}\bm{B}_{t},$
@@
$=\Big{\|}{\int_{0}^{t}\mathbb{E}\left[\nabla_{\theta}U^{\lambda}(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})-\nabla_{\theta}U(\bm{\theta}_{s},\bm{X}_{s})\right]\mathrm{d}s}\Big{\|}^{2}$
$\leq\mathbb{E}\left[\Big{\|}{\int_{0}^{t}[\nabla_{\theta}U^{\lambda}(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})-\nabla_{\theta}U(\bm{\theta}_{s},\bm{X}_{s})]\mathrm{d}s}\Big{\|}^{2}\right],$
@@
$\mathbb{E}\left[\|\bm{X}_{\lambda,t}-\bm{X}_{t}\|^{2}\right]$ $=\mathbb{E}\left[\Big{\|}{\int_{0}^{t}[\nabla_{x}U(\bm{\theta}_{s},\bm{X}_{s})-\nabla_{x}U^{\lambda}(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})]\mathrm{d}s}\Big{\|}^{2}\right].$
@@
$\mathbb{E}\left[\sup_{s\in[0,t]}\|\bm{\theta}_{\lambda,s}-\bm{\theta}_{s}\|^{2}+\mathbb{E}[\|\bm{X}_{\lambda,s}-\bm{X}_{s}\|^{2}]\right]$ $\leq\mathbb{E}\left[\int_{0}^{t}\|\nabla U(\bm{\theta}_{s},\bm{X}_{s})-\nabla U^{\lambda}(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})\|^{2}\mathrm{d}s\right]$
$\leq 2\mathbb{E}\left[\int_{0}^{t}\|\nabla U(\bm{\theta}_{s},\bm{X}_{s})-\nabla U(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})\|^{2}\mathrm{d}s\right]$
$+2\mathbb{E}\left[\int_{0}^{t}\|\nabla U(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})-\nabla U^{\lambda}(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})\|^{2}\mathrm{d}s\right]$
$\leq 2\|\nabla U\|_{\textnormal{Lip}}^{2}\int_{0}^{t}\left(\|\bm{\theta}_{\lambda,s}-\bm{\theta}_{s}\|^{2}+\mathbb{E}[\|\bm{X}_{\lambda,s}-\bm{X}_{s}\|^{2}]\right)\mathrm{d}s$
$+2\mathbb{E}\left[\int_{0}^{t}\|\nabla U(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})-\nabla U^{\lambda}(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})\|^{2}\mathrm{d}s\right].$
@@
$\|\nabla U(v)-\nabla U^{\lambda}(v)\|\leq\|\nabla U\|_{\textnormal{Lip}}(\lambda\|\nabla U(v)\|+\mathcal{O}(\lambda^{2}))\leq\|\nabla U\|_{\textnormal{Lip}}(\lambda\|\nabla U\|_{\textnormal{Lip}}(1+\|v\|)+\mathcal{O}(\lambda^{2})),$ (47)
@@
$\mathbb{E}\left[\int_{0}^{t}\|\nabla U(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})-\nabla U^{\lambda}(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})\|^{2}\mathrm{d}s\right]$ $\leq\lambda^{2}\|\nabla U\|_{\textnormal{Lip}}^{4}\int_{0}^{t}\mathbb{E}\left[1+\|(\bm{\theta}_{\lambda,s},\bm{X}_{\lambda,s})\|^{2}\right]\mathrm{d}s+\|\nabla U\|_{\textnormal{Lip}}^{2}\mathcal{O}(\lambda^{4})t$
@@
Let us denote by $h(t)=\sup_{s\in[0,t]}\|\bm{\theta}_{\lambda,s}-\bm{\theta}_{s}\|^{2}+\mathbb{E}[\|{\bm{X}_{\lambda,s}-\bm{X}_{s}}\|^{2}]$. Then, using the bounds above we have that
@@ D.1.1 Laplace prior with unknown mean $\theta$
We recall that $\phi\in\mathbb{R}^{d}$ is a subdifferential of the $\ell^{1}$-norm at $x\in\mathbb{R}^{d}$ if and only if $\phi_{i}(x)=\text{sign}(x_{i})$ if $x_{i}\neq 0$ and $|\phi_{i}(x)|\leq 1$ otherwise [Parikh and Boyd, 2014].
@@ Results
Algorithm Approx/Iterative Laplace Uniform
NMSE last(%) NMSE avg(%) NMSE last(%) NMSE avg(%)

### v2

@@ Appendix B Theoretical Analysis of Proximal Particle Gradient Descent
where $(\xi_{n})$ for $n\geq 0$ are $d_{x}$-dimensional i.i.d. standard Gaussians.
@@ B.2 Proximal Particle Gradient Descent
We extend the results of Caprio et al. [2024] to provide a nonasymptotic bound for MYPGD. To do so, we consider the following metric on $\mathbb{R}^{d_{\theta}}\times\mathcal{P}_{2}(\mathbb{R}^{d_{x}})$
@@ Appendix C Convergence to Wasserstein gradient flow
We start by proving the following auxiliary result. Let us denote, for any $\theta\in^{d_{\theta}}$ and $\nu\in{\mathcal{P}}(^{d_{x}})$, $g(\theta,\nu):=\int_{{}^{d_{x}}}\nabla_{\theta}U^{\lambda}(\theta,x^{\prime})\nu(x^{\prime})\mathrm{d}x^{\prime}$.
@@
The function $g:^{d_{\theta}}\times{\mathcal{P}}(^{d_{x}})\to\mathbb{R}^{d_{\theta}}$ is Lipschitz continuous in both arguments, i.e.,
@@
$\mathbb{E}\left[\sup_{t\in[0,T]}\left(\|\bm{\theta}_{\lambda,t}-\bm{\theta}^{N}_{t}\|+\|\mathbf{X}_{\lambda,t}-\mathbf{X}_{t}^{j,N}\|\right)\right]\leq\frac{\sqrt{2}(\sqrt{C_{T}}\lambda^{-1}+\sqrt{T})e^{2T\lambda^{-1}}}{N^{1/2}}$ (48)
where $C_{T}:=\sup_{t\leq T}\mathbb{E}\left[|\bm{\theta}^{N}_{t}|^{2}+|\mathbf{X}_{t}^{j,N}|^{2}\right]<\infty$, for any $j=1,\dots,N$.
@@
$\mathbb{E}\left[\sup_{t\in[0,T]}\left(\|\bm{\theta}_{\lambda,t}-\bm{\theta}_{t}\|^{2}+\|\mathbf{X}_{\lambda,t}-\mathbf{X}_{t}\|^{2}\right)\right]\leq\left(\lambda^{2}\|\nabla U\|_{\textnormal{Lip}}^{4}C_{T}+\|\nabla U\|_{\textnormal{Lip}}^{2}\mathcal{O}(\lambda^{4})\right)T\exp(2\|\nabla U\|_{\textnormal{Lip}}^{2}T),$
where $C_{T}:=\sup_{t\leq T}\mathbb{E}\left[|\bm{\theta}_{\lambda,t}|^{2}+|\mathbf{X}_{\lambda,t}|^{2}\right]<\infty$. It follows that, as $\lambda\to 0$, (47) converges to (49) in $\mathbb{L}^{2}$.
@@
$\mathbf{X}_{t}$ $=X_{0}-\int_{0}^{t}\nabla_{x}U(\bm{\theta}_{s},\mathbf{X}_{s})\mathrm{d}s+\sqrt{2}\bm{B}_{t},$
@@
$=\Big{\|}{\int_{0}^{t}\mathbb{E}\left[\nabla_{\theta}U^{\lambda}(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})-\nabla_{\theta}U(\bm{\theta}_{s},\mathbf{X}_{s})\right]\mathrm{d}s}\Big{\|}^{2}$
$\leq\mathbb{E}\left[\Big{\|}{\int_{0}^{t}[\nabla_{\theta}U^{\lambda}(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})-\nabla_{\theta}U(\bm{\theta}_{s},\mathbf{X}_{s})]\mathrm{d}s}\Big{\|}^{2}\right],$
@@
$\mathbb{E}\left[\|\mathbf{X}_{\lambda,t}-\mathbf{X}_{t}\|^{2}\right]$ $=\mathbb{E}\left[\Big{\|}{\int_{0}^{t}[\nabla_{x}U(\bm{\theta}_{s},\mathbf{X}_{s})-\nabla_{x}U^{\lambda}(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})]\mathrm{d}s}\Big{\|}^{2}\right].$
@@
$\mathbb{E}\bigg{[}\sup_{s\in[0,t]}\|\bm{\theta}_{\lambda,s}$ $-\bm{\theta}_{s}\|^{2}+\mathbb{E}[\|\mathbf{X}_{\lambda,s}-\mathbf{X}_{s}\|^{2}]\bigg{]}\leq\mathbb{E}\left[\int_{0}^{t}\|\nabla U(\bm{\theta}_{s},\mathbf{X}_{s})-\nabla U^{\lambda}(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})\|^{2}\mathrm{d}s\right]$
$\leq 2\mathbb{E}\left[\int_{0}^{t}\|\nabla U(\bm{\theta}_{s},\mathbf{X}_{s})-\nabla U(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})\|^{2}\mathrm{d}s\right]$
$+2\mathbb{E}\left[\int_{0}^{t}\|\nabla U(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})-\nabla U^{\lambda}(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})\|^{2}\mathrm{d}s\right]$
$\leq 2\|\nabla U\|_{\textnormal{Lip}}^{2}\int_{0}^{t}\left(\|\bm{\theta}_{\lambda,s}-\bm{\theta}_{s}\|^{2}+\mathbb{E}[\|\mathbf{X}_{\lambda,s}-\mathbf{X}_{s}\|^{2}]\right)\mathrm{d}s$
$+2\mathbb{E}\left[\int_{0}^{t}\|\nabla U(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})-\nabla U^{\lambda}(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})\|^{2}\mathrm{d}s\right].$
@@
$\|\nabla U(v)-\nabla U^{\lambda}(v)\|$ $\leq\|\nabla U\|_{\textnormal{Lip}}(\lambda\|\nabla U(v)\|+\mathcal{O}(\lambda^{2}))$
$\leq\|\nabla U\|_{\textnormal{Lip}}(\lambda\|\nabla U\|_{\textnormal{Lip}}(1+\|v\|)+\mathcal{O}(\lambda^{2})),$ (50)
@@
$\mathbb{E}$ $\left[\int_{0}^{t}\|\nabla U(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})-\nabla U^{\lambda}(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})\|^{2}\mathrm{d}s\right]$
$\leq\lambda^{2}\|\nabla U\|_{\textnormal{Lip}}^{4}\int_{0}^{t}\mathbb{E}\left[1+\|(\bm{\theta}_{\lambda,s},\mathbf{X}_{\lambda,s})\|^{2}\right]\mathrm{d}s+\|\nabla U\|_{\textnormal{Lip}}^{2}\mathcal{O}(\lambda^{4})t$
@@
Let us denote by $h(t)=\sup_{s\in[0,t]}\|\bm{\theta}_{\lambda,s}-\bm{\theta}_{s}\|^{2}+\mathbb{E}[\|{\mathbf{X}_{\lambda,s}-\mathbf{X}_{s}}\|^{2}]$. Then, using the bounds above we have that
@@ D.1.1 Laplace prior with unknown mean $\theta$
We recall that $\phi\in^{d}$ is a subdifferential of the $\ell^{1}$-norm at $x\in^{d}$ if and only if $\phi_{i}(x)=\text{sign}(x_{i})$ if $x_{i}\neq 0$ and $|\phi_{i}(x)|\leq 1$ otherwise [Parikh and Boyd, 2014].
@@ Results
Algorithm Approx/ Laplace Uniform
Iterative NMSE last(%) NMSE avg(%) NMSE last(%) NMSE avg(%)